OpenStudy (anonymous):

Show each step of your work. four times the quantity thirty-two minus two end quantity plus twelve divided by four times three

OpenStudy (anonymous):

= 4 * (32 -2) + 12/4*3 = 4 * 30 + 9 = 120 + 9 = 129